You have a project to plant corn in 4 fields each having an area of 20 ha (200000 m2). Each field is to take one day to plant, and $1,000 has been budgeted per field. The fields were planned to be planted one after the other. Today is the end of day 3. Using the following project status chart, calculate PV, EV, etc., in the spaces provided.

Activity Day1 Day 2 Day3 Day 4 Status at the end of day 3
Field 1 S-------F Complete, spent $1000
Field 2 S------PF -------F Complete spent $1200
Field 3 PS--S--PF 50% done spent $600
Field 4 PS--PF Not Yet Started

Key
S Start
PS planned start
F Finish
PF Planned Finish

Calculate the following and give an interpretation for each answer:

PV

EV

AC

BAC

CV

CPI

SV

SPI

EAC

ETC
Answer #1

BudgetAtCompletion = 4Days for-4 fields* 1000 /day= 4000

Project was Budgeted at $4000 as the BAC is $4000

Since by the end of Day 3, 75% of the Work should be completed as per Plan

PV = (Planned Work * BAC)=0.75 * $4000 = $3000

EV= (Actual Work * BAC) = 0.625 * 4000 = $2500

AC= 1000+1200+600 = $2800

CV= EV-AC=2500-2800 = -300

A negative CV implies that we have over spent compared to the budget

CPI=\frac{BAC}{AC} =\frac{4000}{2800} = 1.428

A Higher than 1 CPI implies the project is so far within the Budget

SV = EV-PV =2500-3000= -500

A Negative SV implies we are behind schedule

SPI =\frac{EV}{PV} =\frac{2500}{3000} = 0.833

SPI of 0.83 implies we have Lost 17% of the available time

EAC= AC + (BAC-EV)= 2800 + (4000-2500)= $4300

If the project continues as per plan apart from the Budgeted overshoot without any further problems we may end up wth $4300 at the end of project against the Budgeted $4000

ETC = EAC - AC = $4300 - $2800 = $1500

To Complete the proect we will have to spend further $1500